OpenStudy (anonymous):

MEDAL AND FAN! A system of equations is shown below. -3x + 7y = -16 -9x + 5y = 16 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. (6 points) Part B: Show that the equivalent system has the same solution as the original system of equations. (4 points)

OpenStudy (anonymous):

Replace (2) by adding (2) to two times (1): 2 times (1): -6x + 14y = -32 Add (2): -9x + 5y = 16 -------------------- -15x + 19y = -16 So the second pair of equations are: -3x + 7y = -16 ---- (1) -15x + 19y = -16 ----- (3)

OpenStudy (anonymous):

@SolomonZelman can you help me with part B please?

OpenStudy (anonymous):

@Michele_Laino Can you please help me?

OpenStudy (anonymous):

@Secret-Ninja can you help me with part B?

OpenStudy (michele_laino):

ok! please now you have to solve the second part

OpenStudy (anonymous):

how do I do that?

OpenStudy (michele_laino):

another solution, can be this: I multiply your first equation by (-3) and then I add it to the second equation, namely: (-3) times (1)--->9x-21y=48 add (2)-------> -9x+5y=16 ------------------ -16y=64 SO the second pair of equations are: -3x+7y=-16 -16y=64

OpenStudy (anonymous):

Wait so my part A wasn't right?

OpenStudy (michele_laino):

no, no, it's right!

OpenStudy (anonymous):

Oh :o so what did u just do right there? Was that a different solution to part A?

OpenStudy (michele_laino):

what I want you to understand is that the purpose of manipulating the equations of a system is to obtain the simplified equations

OpenStudy (anonymous):

Okays I understand how manipulating equations of a system can obtain simplified equations but how do I apply that to part b?

OpenStudy (michele_laino):

for example, both results, namely mine and yours are correct, even if, by my answer we have the second equation already simplified, and we can solve it rapidly

OpenStudy (michele_laino):

solution of second system is: x=-4, y=-4 do you agree?

OpenStudy (anonymous):

No the second system is -3x+7y=-16 and -15x+19y=-16

OpenStudy (anonymous):

where did you get x=-4 and y=-4?

OpenStudy (michele_laino):

please, I refer to my second system!

OpenStudy (anonymous):

Oh! Your second system

OpenStudy (anonymous):

:D I get it! Since you simplified 64 and -16

OpenStudy (michele_laino):

that's right!

OpenStudy (michele_laino):

now you have to solve your first system, and then discover that the resultant solution is equal to that of the second system

OpenStudy (anonymous):

I know I need to Solve (1) and (2) for x,y and Solve (1) and (3) for x,y

OpenStudy (michele_laino):

I solve your first system: from first equation, we have: \[x=\frac{ 7y +16}{ 3}\] inserting that expression, into your second equation, we have: \[-9*\frac{ 7y+16 }{ 3 }+5y=16\] I simplify: \[-16y=64\] from which y=-4 Now I substitute y=-4 into the expression for x, namely: \[x=\frac{ 7y+16 }{ 3 }=\frac{ 7*(-4)+16 }{ 3 }=\frac{ -12 }{ 3 }=-4\] then solution of your first system is x=-4, y=-4 which is equal to solution of my second system. In other words your first system and my second system are equivalent

OpenStudy (anonymous):

Thank you @Michele_Laino For all your help ^_^

OpenStudy (michele_laino):

Thank you @Sandybottoms1432